# Numerical model for pellet rocket acceleration in PELOTON

### v1

@@ Abstract
keywords:
pellet ablation, pellet rocket acceleration, shattered pellet injection, pellet fueling, JET
articletype: Paper
@@ 2.1 Pellet surface ablation and low magnetic Reynolds number MHD approximation for ablation cloud
where $d/dt$ is the Lagrangian time derivative, $\mathbf{u}$, $\rho$ and $\epsilon$ are the velocity, density and specific internal energy, respectively, $P$ is the pressure, $\mathbf{B}=(0,0,B_{||})$ is the prescribed magnetic field induction, $\mathbf{J}$ is the current density computed as $\mathbf{J}=\sigma_{\perp}\mathbf{u}\times\mathbf{B}$, and $\sigma_{\perp}$ is the transverse electric conductivity. Last term in (6), the energy deposition of hot plasma electrons streaming into the ablation cloud, is described in Section 2.2. Finally, the ablation cloud experiences grad-$\mathbf{B}$ drift in the transverse direction to the magnetic field represented by the term $\mathbf{f}_{D}$ in (5). In present work, we use a simplified grad-$\mathbf{B}$ drift model of [3] as we are interested in the state of ablation cloud sufficiently close to the pellet in transverse direction to the magnetic field. In particular, the transverse drift acceleration $dv_{\mathrm{D}}/dt$ in the LFS direction is governed by the following equation
@@
Here $\langle A\rangle$ is the integral of quantity $A$ along a magnetic field line, $\langle A\rangle\equiv\int_{-L}^{L}A\mathrm{~d}z$, where $L$ is the cloud length, $R$ denotes the magnetic field radius of curvature, $P_{\infty}$ is the ambient plasma pressure, and $\left\langle P-P_{\infty}\right\rangle$ is called the drive integral. We use a more advanced model for cross-field transport of [10] if modeling of the plasmoid drift over larger distances is necessary.
@@ 2.2 Electron heat deposition
$\nabla\cdot\mathbf{q}=\frac{q_{\infty}n_{\rm e}(r,z)}{\tau_{\rm eff}}\left[g\left(u_{+}\right)+g\left(u_{-}\right)\right],$ (9)
@@
$g(u)=u^{\frac{1}{2}}K_{1}\left(u^{\frac{1}{2}}\right)/4$, where $K_{1}$ is the standard modified Bessel function of the second kind and $T_{\mathrm{e}\infty}$ is the temperature of the plasma electrons. Quantities $u_{\pm}=\frac{\tau_{\pm}}{\tau_{\mathrm{cff}}}$ are dimensionless opacities of the cloud with respect to each (+ or -) longitudinal direction from a point of interest, where the respective line integrated densities of the ablation electrons (bound and free) are,
@@
where $e$ is the elementary charge of electron. The Coulomb logarithm $\ln\Lambda$ here pertains to inelastic scattering of fast electrons off atomic (bound) electrons in the neutral gas target,
$\ln\Lambda=\ln\left(\frac{E}{I^{*}}\sqrt{\frac{e}{2}}\right)$ (15)
where $e$ is Napier’s constant (not to be confused with elementary charge $e$ in the equation for $\tau_{\infty}$ ), and $I^{*}$ is the mean excitation energy for neutral atoms. The Coulomb logarithm is evaluated at energy $E\approx 2T_{\mathrm{e}\infty}$ since that is the average energy per particle in the cloud from the distribution of semi-isotropic incident Maxwellian electrons.
@@
where the attenuation coefficient $\mu$ is a function of the normalized cloud opacity $u=u_{+}+u_{-}$
$\mu(u)=\left[1+\sqrt{\frac{u}{2}}-\frac{u}{4}+\frac{1}{2}\left(\frac{u}{2}\right)^{\frac{3}{2}}\right]\exp\left(-\sqrt{\frac{u}{2}}\right)-\left(\frac{u^{2}}{8}\right)E_{1}(\sqrt{\frac{u}{2}}).$ (17)
@@
$q_{\pm}=q_{\infty}\frac{1}{2}u_{\pm}K_{2}\left(u_{\pm}^{\frac{1}{2}}\right).$ (18)
@@ 2.3 Equation of state for deuterium-neon mixture
Let $C$ denote the ratio of number of deuterium to neon nuclei. We wish to solve for the particle concentrations $f_{0}^{N},f_{1}^{N},\ldots,f_{10}^{N},f_{D},f_{D+}$, and $f_{e}$, where $f_{i}^{N}$ is the concentration of the i-th ionized Neon ion, $f_{D}$ is the concentration of deuterium atoms, $f_{D^{+}}$ is the concentration of deuterium ions, and $f_{e}$ is the concentration of electrons. These fractions will be used to compute quantities such as temperature and sound speed. We setup a nonlinear system of equations based on 10 Saha equations for neon, 2 Saha equations for deuterium, along with conservation of mass and conservation of charge equations. These equations are as follows:
$\frac{f_{m+1}^{N}f_{e}}{f_{m}^{N}}=\frac{2}{N_{t}}\frac{u_{m+1}^{N}}{u_{m}^{N}}\left(\frac{2\pi m_{e}kT}{h^{2}}\right)^{\frac{3}{2}}\exp\left(-\frac{I_{m+1}^{N}}{kT}\right),\quad m=0,1,\ldots,9$ (19)
@@
$f_{0}^{N}+f_{1}^{N}+f_{2}^{N}+\ldots+f_{10}^{N}=1$ (22)
$f_{e}=\frac{f_{1}^{N}}{1+C}+\frac{2f_{2}^{N}}{1+C}+\ldots+\frac{10f_{10}^{N}}{1+C}+\frac{f_{D^{+}}}{1+1/C},$ (23)
where $h$ is the Planck constant, $k$ is the Boltzmann constant, $m_{e}$ is the electron mass, $u_{m}$ are known electron partition functions, and $I_{m}$ are the successive ionization potentials. For deuterium, the ionization energy is $\epsilon_{i}=13.6\mathrm{eV}$ and the dissociation energy is $\epsilon_{d}=4.48\mathrm{eV}$, and finally we have $\beta_{i}=\frac{3}{2}$ and the parameter $\beta_{d}=0.327$.
@@ 2.4 Plasma Cooling Module
PELOTON’s Cooling Module implements a physics model that discretizes the plasma into flux volumes, distributes ablated material within the corresponding flux volumes and accounts for ionization and other energy losses including radiation, Ohmic heating by fixed toroidal currents, and the energy exchange between ions and electrons. In simulation results performed by the cooling module presented herein, radiation modeling is not applied. Plasma profiles predicted by PELOTON’s cooling model have been compared with simulations from two other plasma evolution codes, JOREK and INDEX [17, 18].
PELOTON’s Cooling Module represents the plasma using a number of discrete flux volumes, each of which has a spatially constant thermodynamic state. Consider an ablating SPI fragment moving with a certain velocity in a tokamak. During an interval of time $dt$, chosen sufficiently small so that the background plasma temperature and density can be considered constant during the interval, the pellet ablates material that increases the ion and electron densities by $dn_{i}$ and $dn_{e}$, respectively. These quantities are computed as:
$dn_{i}=\frac{G\,dt}{V\,m_{i}},\quad dn_{e}=f_{e}dn_{i},$ (26)
@@
$\frac{dP_{e}}{dt}=-Q_{ion}-Q_{rad}+n_{e}\nu_{ie}(T_{i}-T_{e})+\frac{\gamma-1}{R^{2}}\eta(T_{e})j^{2}$ (27)
@@
where $Q_{ion}$ is the ionization power density and $Q_{rad}$ is the radiation power density (for neon or mixed pellets). The last term in the right-hand side of equation (27) describes the Ohmic heating computed using the equilibrium current density profile, where the resistivity $\eta$ is computed using expressions from the NRL Plasma Formulary [19]. The last term in the right-hand side of equation (28) describes the ion-electron collisional energy exchange and $\nu_{ie}$ is the electron–ion collision rate. We assume an ideal gas equation of state and find new background temperatures as
@@ 4.1 Simulations at conditions of JET SPI and scaling law
We used 20 simulation data points to determine this fit which covered the approximate parameter ranges $T_{e}\in$ [450 eV, 2000 eV], $n_{e}\in$ [6e19 $m^{-}3$, 3e20 $m^{-}3$], $r_{p}\in$ [0.2 mm, 1.8 mm], where $r_{p}$ is the pellet radius. Comparison of this fit and data points computed by PELOTON gives the mean absolute relative error of approximately 13 percent. This error is mainly due to numerical noise in computed rocket acceleration values because the mean signed relative error is only one percent. Therefore, we believe that (30) accurately represents $\Delta P$ values for the given parameter range. We observed that the rocket acceleration related pressure difference is practically independent on the pellet radius: the regression fit into numerical data yielded the following relation
@@ 4.2 Validation with deuterium SPI experiment in JET
During each fragment ablation, the ablated plasmoids expand longitudinally along magnetic field lines, filling the flux volumes, and experience drift in the $\nabla R$ direction. Due to plasmoid drift, the location of effective material deposition influencing the ambient plasma states is shifted with respect to the plasmoid location. This significantly changes plasma profiles and the resulting ablation dynamics of fragments. A fixed 40 cm shift was used in [18] to match experimental measurements. The actual plasmoid drift, however, depends on the states of plasmoids and the ambient plasma [25, 26]. In this work, we adopt a variable plasmoid shift ad hoc model approximating [26]. Assuming that the plasmoid drift is generally proportional to fragment’s ablation rate, we apply $R_{shift}=\alpha*G$, where G is the ablation rate and $\alpha$ is an adjustable parameter. We use a value of $\alpha=20\frac{m*s}{kg}$ as it results in a reasonable agreement in the overall penetration depth of SPI fragments with experiments. In addition, the energy deposition accounts for the energy gained by ablated material due to dissociation, ionization, and heating while drifting through different parts of the plasma. This energy (100 eV per ablated atom) is deducted from the location of ablating fragment and is added at the deposition location.
Figure 3(b)(a) shows the evolution of fragment radii. The overall ablation process lasts 4 ms. Figure 3(b)(b) shows drift induced shifts experiencing by the corresponding plasmoids. Colors in Figure 3(b) are consistent among three plots showing, for example, that the largest fragment experienced the shift of 0.8 m at the peak of its ablation rate. The evolution of rocket acceleration values experienced by all fragments are depicted in 3(b)(c). As expected, the rocket acceleration rapidly increased during short intervals of time at the end of each fragment lifetime due to decreasing mass, and pressure difference independent of fragment radius. Rocket acceleration in this simulation reached maximum values greater than $5\times 10^{6}\,m/s^{2}$. Typical rocket acceleration values remained below $10^{6}\,m/s^{2}$ for most of fragment lifetimes.
@@
Finally, we discuss the plasma profiles obtained by PELOTON’s dilution cooling code. Plasma profiles in JET at selected times are displayed in figure 3(d). For selected SPI fragment trajectories shown in Figure 3(c), electron temperature and density experienced by each fragment as a function of time is shown in figure 3(e).
(d) (a) Electron temperature and (b) electron density predicted by PELOTON’s cooling module at selected times.
(e) (a) Temperature experienced by selected fragments depicted in Figure 3(c) (b) as a function of time. b) Density experienced by selected fragments depicted in Figure 3(c) (b) as a function of time.
Comparisons with JOREK simulations of JET SPI show that PELOTON’s cooling code over-predicts density at the plasma edge when the shift model for the ablated material deposition of [18] is applied to model the effect of the plasmoid drift. PELOTON’s cooling module does not include diffusion across flux surfaces and MHD effects. We observe density gradients near the plasma edge caused by shifted deposition of material and we attribute this to the missing physics effects that would diffuse plasma particles across the scrape-off-layer (SOL). We performed tests designed to explore the impact of lower densities at the plasma edge and corresponding higher temperatures on SPI fragment trajectories. We artificially reduced densities when specified cut-off levels are reached and correspondingly increased temperatures. We did not display results for conciseness, but observed shorter fragment lifetimes, particularly among smaller fragments. The overall qualitative behavior of the plume remained very similar.
4.2.1 Influence of neon component

### v2

@@ Abstract
Keywords: pellet ablation, pellet rocket acceleration, shattered pellet injection, pellet fueling, JET
@@ 2.1 Pellet surface ablation and low magnetic Reynolds number MHD approximation for ablation cloud
where $d/dt$ is the Lagrangian time derivative, $\mathbf{u}$, $\rho$ and $\epsilon$ are the velocity, density and specific internal energy, respectively, $P$ is the pressure, $\mathbf{B}=(0,0,B_{||})$ is the prescribed magnetic field induction, $\mathbf{J}$ is the current density computed as $\mathbf{J}=\sigma_{\perp}\mathbf{u}\times\mathbf{B}$, and $\sigma_{\perp}$ is the transverse electric conductivity. Last term in (6), the energy deposition of hot plasma electrons streaming into the ablation cloud, is described in Section 2.2. Finally, the ablation cloud experiences grad-$\mathbf{B}$ drift in the transverse direction to the magnetic field governed by force $\mathbf{f}_{D}$ in (5). In present work, we use a simplified grad-$\mathbf{B}$ drift model as we are interested in the state of ablation cloud sufficiently close to the pellet in transverse direction to the magnetic field. In particular, the transverse drift acceleration $dv_{\mathrm{D}}/dt=\mathbf{f}_{D}/\rho$ in the LFS direction is governed by the following equation
@@
Here $\langle A\rangle$ is the integral of quantity $A$ along a magnetic field line, $\langle A\rangle\equiv\int_{-L}^{L}A\mathrm{~d}z$, where $L$ is the cloud length, $R$ denotes the magnetic field radius of curvature, $P_{\infty}$ is the ambient plasma pressure, and $\left\langle P-P_{\infty}\right\rangle$ is called the drive integral. We use a more advanced model for cross-field transport of [10] if modeling of the plasmoid drift over larger distances is necessary. The Alfven wave drag and the external connection current drag terms considered in [10] are important at larger distances from the pellet in the transverse direction compared to length scales for computing ablation rates and rocket acceleration values.
@@ 2.2 Electron heat deposition
$\nabla\cdot\mathbf{q}=\frac{q_{\infty}n_{\rm e}(r,z)}{\tau_{\rm eff}}\left[g\left(w_{+}\right)+g\left(w_{-}\right)\right],$ (9)
@@
$g(w)=w^{\frac{1}{2}}K_{1}\left(w^{\frac{1}{2}}\right)/4$, where $K_{1}$ is the standard modified Bessel function of the second kind and $T_{\mathrm{e}\infty}$ is the temperature of the plasma electrons. Quantities $w_{\pm}=\frac{\tau_{\pm}}{\tau_{\mathrm{cff}}}$ are dimensionless opacities of the cloud with respect to each (+ or -) longitudinal direction from a point of interest, where the respective line integrated densities of the ablation electrons (bound and free) are,
@@
where $e$ is the elementary charge of electron and $Z_{*}$ denotes the average ionization state of partially ionized plasma. The Coulomb logarithm $\ln\Lambda$ here pertains to inelastic scattering of fast electrons off atomic (bound) electrons in the neutral gas target,
$\ln\Lambda=\ln\left(\frac{E}{I^{*}}\sqrt{\frac{e(=\exp(1))}{2}}\right),$ (15)
where $I^{*}$ is the mean excitation energy for neutral atoms. The Coulomb logarithm is evaluated at energy $E\approx 2T_{\mathrm{e}\infty}$ since that is the average energy per particle in the cloud from the distribution of semi-isotropic incident Maxwellian electrons.
@@
where $\rm erfc$ is the Gauss error function and the attenuation coefficient $\mu$ is a function of the normalized cloud opacity $w=w_{+}+w_{-}$
$\mu(w)=\left[1+\sqrt{\frac{w}{2}}-\frac{w}{4}+\frac{1}{2}\left(\frac{w}{2}\right)^{\frac{3}{2}}\right]\exp\left(-\sqrt{\frac{w}{2}}\right)-\left(\frac{w^{2}}{8}\right)E_{1}(\sqrt{\frac{w}{2}}),$ (17)
where $E_{1}$ is the exponential integral.
@@
$q_{\pm}=q_{\infty}\frac{1}{2}w_{\pm}K_{2}\left(w_{\pm}^{\frac{1}{2}}\right),$ (18)
where $q_{\pm}$ is either $q_{+}$ or $q_{-}$, the electron heat flux impacting the left or right sides of the pellet, correspondingly.
@@ 2.3 Equation of state for deuterium-neon mixture
Let $C$ denote the ratio of number of deuterium to neon nuclei. We wish to solve for the particle fractions $f^{0}_{Ne},f^{1}_{Ne},\ldots,f^{10}_{Ne},f_{D},f_{D+}$, and $f_{e}$, where $f^{i}_{Ne}$ is the fraction of the i-th ionized Neon ion, $f_{D}$ is the fraction of deuterium atoms, $f_{D^{+}}$ is the fraction of deuterium ions, and $f_{e}$ is the fraction of electrons. These fractions are used to compute quantities such as temperature, sound speed, and electric conductivity. We setup a nonlinear system of equations based on 10 Saha equations for neon, 2 Saha equations for deuterium, along with conservation of mass and conservation of charge equations. These equations are as follows:
$\frac{f^{m+1}_{Ne}f_{e}}{f^{m}_{Ne}}=\frac{2}{N_{t}}\frac{g_{m+1}^{N}}{g_{m}^{N}}\left(\frac{2\pi m_{e}kT}{h^{2}}\right)^{\frac{3}{2}}\exp\left(-\frac{I_{m+1}^{N}}{kT}\right),\quad m=0,1,\ldots,9$ (19)
@@
$f^{0}_{Ne}+f^{1}_{Ne}+f^{2}_{Ne}+\ldots+f^{10}_{Ne}=1$ (22)
$f_{e}=\frac{f^{1}_{Ne}}{1+C}+\frac{2f^{2}_{Ne}}{1+C}+\ldots+\frac{10f^{10}_{Ne}}{1+C}+\frac{f_{D^{+}}}{1+1/C},$ (23)
where $h$ is the Planck constant, $k$ is the Boltzmann constant, $m_{e}$ is the electron mass, $g_{m}$ are known electron partition functions, and $I_{m}$ are the successive ionization potentials. For deuterium, the ionization energy is $\epsilon_{i}=13.6\mathrm{eV}$ and the dissociation energy is $\epsilon_{d}=4.48\mathrm{eV}$, and finally we have $\beta_{i}=\frac{3}{2}$ and the parameter $\beta_{d}=0.327$.
@@ 2.4 Plasma Cooling Module
PELOTON’s Cooling Module implements a physics model that discretizes the plasma into flux volumes, distributes ablated material within the corresponding flux volumes and accounts for ionization and other energy losses including radiation, Ohmic heating by fixed toroidal currents, and the energy exchange between ions and electrons. In simulation results performed by the cooling module presented herein, radiation modeling is not applied. Plasma profiles predicted by PELOTON’s cooling model have been compared with simulations from two other plasma evolution codes, JOREK and INDEX [17, 18]. PELOTON’s Cooling Module differs from INDEX by using a temporally fixed flux surface geometry and current profile. Additionally, transport and radiation models are not applied in this study. PELOTON’s Cooling Module incorporates pellet ablation and rocket acceleration models referencing data from PELOTON simulations.
PELOTON’s Cooling Module represents the plasma using a number of discrete flux volumes, each of which has a spatially constant thermodynamic state. Consider an ablating SPI fragment moving with a certain velocity in a tokamak. During an interval of time $dt$, chosen sufficiently small so that the background plasma temperature and density can be considered constant during the interval, the pellet ablates material that increases the ion and electron densities by $dn_{i}$ and $dn_{e}$, respectively. These quantities are computed by solving the following equations in the time interval $dt$:
$\frac{dn_{i}}{dt}=\frac{G}{V\,m_{i}},\quad\frac{dn_{e}}{dt}=f_{e}\frac{dn_{i}}{dt},$} (26)
@@
$\frac{dP_{e}}{dt}=-Q_{{sink}}-Q_{rad}+n_{e}\nu_{ie}(T_{i}-T_{e})+\frac{\gamma-1}{R^{2}}\eta(T_{e})j^{2}$ (27)
@@
where $Q_{{sink}}$ is the power density loss due to ionization, dissociation, and if plasmoid drift is accounted for, cloud heating by plasma electrons is included in this term. $Q_{rad}$ is the radiation power density (for neon or mixed pellets). The last term in the right-hand side of equation (27) describes the Ohmic heating computed using the equilibrium current density profile, where the resistivity $\eta$ is computed using expressions from the NRL Plasma Formulary [19]. The last term in the right-hand side of equation (28) describes the ion-electron collisional energy exchange and $\nu_{ie}$ is the electron–ion collision rate. We assume an ideal gas equation of state and find new background temperatures as
@@ 4.1 Simulations at conditions of JET SPI and scaling law
We used 20 simulation data points to determine this fit which covered the approximate parameter ranges $T_{e}\in$ [450 eV, 2000 eV], $n_{e}\in$ [6e19 $m^{-}3$, 3e20 $m^{-}3$], $r_{p}\in$ [0.2 mm, 1.8 mm], where $r_{p}$ is the pellet radius. Exact simulation parameters and results are detailed in table 4 within the appendix. Comparison of this fit and data points computed by PELOTON gives the mean absolute relative error of approximately 13 percent. This error is mainly due to numerical noise in computed rocket acceleration values because the mean signed relative error is only one percent. Therefore, we believe that (30) accurately represents $\Delta P$ values for the given parameter range. We observed that the rocket acceleration related pressure difference is practically independent on the pellet radius: the regression fit into numerical data yielded the following relation
@@ 4.2 Validation with deuterium SPI experiment in JET
During each fragment ablation, the ablated plasmoids expand longitudinally along magnetic field lines, filling the flux volumes, and experience drift in the $\nabla R$ direction. Due to plasmoid drift, the location of effective material deposition influencing the ambient plasma states is shifted with respect to the plasmoid location. This significantly changes plasma profiles and the resulting ablation dynamics of fragments. We adopt a ‘teleportation’ model similar to [18], whereby the effects of plasmoid drift are approximated by depositing material locally around a location shifted in the direction of the plasmoid drift from the ablation location. We adopt this model as an approximation and do not explicitly track the material deposition during the plasmoid drift process. A fixed 40 cm shift was used in [18] to match experimental measurements. The actual plasmoid drift, however, depends on the states of plasmoids and the ambient plasma [25, 26]. In this work, we adopt a variable plasmoid shift ad hoc model approximating [26]. Assuming that the plasmoid drift is generally proportional to fragment’s ablation rate, we apply $R_{shift}=\alpha*G$, where G is the ablation rate and $\alpha$ is an adjustable parameter. We use a value of $\alpha=20\frac{m*s}{kg}$ as it results in a reasonable agreement in the overall penetration depth of SPI fragments with experiments, as fragment trajectories reach comparable depth as the curve indicative of the cloud edge in terms of the flux coordinate . In addition, the energy deposition accounts for the energy gained by ablated material due to dissociation, ionization, and heating while drifting through different parts of the plasma. We adopt the approximation that all such energy is provided to the ablation cloud by the region local to the ablating fragment and do not explicitly track energy gain during the drift process. This energy (100 eV per ablated atom) is deducted from the location of ablating fragment and is added at the deposition location.
Figure 3(b)(a) shows the evolution of fragment radii. The overall ablation process lasts 4 ms. Figure 3(b)(b) shows drift induced shifts experiencing by the corresponding plasmoids. Colors in Figure 3(b) are consistent among three plots showing, for example, that the largest fragment experienced the shift of 0.8 m at the peak of its ablation rate. In such cases, the plasmoid drifts into the SOL. From the perspective of the material deposition in the plasma, such material and the energy gained during drift is not added to the plasma and is effectively discarded from the simulation. The evolution of rocket acceleration values experienced by all fragments are depicted in 3(b)(c). As expected, the rocket acceleration rapidly increased during short intervals of time at the end of each fragment lifetime due to decreasing mass, and pressure difference independent of fragment radius. Rocket acceleration in this simulation reached maximum values greater than $5\times 10^{6}\,m/s^{2}$. Typical rocket acceleration values remained below $10^{6}\,m/s^{2}$ for most of fragment lifetimes.
@@
To clarify the influence of various factors on the rocket acceleration, we plotted the electrostatic shielding potential across the ablated cloud and the corresponding reduction of the incoming hot electron density $\exp(-\Phi)$ (blue and red lines in Figure 3, correspondingly). As expected, the potential is steeper on the HFS, providing less shielding at the same distance from the pellet compared to the LFS. However, the electrostatic shielding potential effect on the rocket acceleration is much smaller compared to the difference of decay of hot electrons on both sides of the pellet due to the ablation cloud asymmetry. We estimated that the energy deposition per unit mass into ablated plasma at 1 mm off the pellet surface was about 35% higher on the HFS compared to the LFS. The corresponding difference in the incoming hot electron reduction $\exp(-\Phi)$ was only 1.4%. We also verified that the low magnetic Reynolds approximation is accurate for the denser part of the plasmoid defining the ablation rate and rocket acceleration.
Figure 3: Electrostatic shielding potential across the ablation cloud in the direction of grad-B drift (blue line) and reduction of the incoming hot electron density by electrostatic shielding (red line)
Finally, we discuss the plasma profiles obtained by PELOTON’s dilution cooling code. Plasma profiles in JET at selected times are displayed in Figure 5(a). We observe the increase of plasma density and reduction of temperature with the deposition of new ablated material. For selected SPI fragment trajectories shown in Figure 3(c)(b), electron temperature and density experienced by each fragment as a function of time is shown in figure 5(b). Most of selected fragments ablated at low plasma $T_{e}$ and high $n_{e}$. Despite reduced values of their rocket acceleration compared to the fragment that experienced plasma $T_{e}$ above 2 keV (purple line), trajectories of these fragments deviated more from the straight line because of their long lifetime / duration of the rocket acceleration. The exception was the fragment with shortest penetration depth due to small initial size, plotted with a yellow line (see Figure 3(c)(b)).
(a) (a) Electron temperature and (b) electron density predicted by PELOTON’s cooling module at selected times.
(b) (a) Temperature experienced by selected fragments depicted in Figure 3(c) (b) as a function of time. b) Density experienced by selected fragments depicted in Figure 3(c) (b) as a function of time.
PELOTON’s cooling code over-predicts density at the plasma edge when the shift model for the ablated material deposition of [18] is applied to model the effect of the plasmoid drift. PELOTON’s cooling module does not include diffusion across flux surfaces and MHD effects. We observe density gradients near the plasma edge caused by shifted deposition of material and we attribute this to the missing physics effects that would diffuse plasma particles across the scrape-off-layer (SOL). We performed tests designed to explore the impact of lower densities at the plasma edge and corresponding higher temperatures on SPI fragment trajectories. We artificially reduced densities when specified cut-off levels are reached and correspondingly increased temperatures. We did not display results for conciseness, but observed shorter fragment lifetimes, particularly among smaller fragments. The overall qualitative behavior of the plume remained very similar. The overall behavior is similar because of the shifted deposition of ablated material. The over-predicted density buildup at the plasma edge mostly impacts the relatively small fragments in the trailing end of the SPI plume.
4.3 Influence of neon component